OpenStudy (vanz):

How do you find the value of K so each given trinomial is a perfect square? x^2+12x+k

OpenStudy (anonymous):

36

OpenStudy (vanz):

how do you get that? :O

OpenStudy (anonymous):

12/2 = 6 6^2 = 36

OpenStudy (vanz):

could i do the same with this 9x^2-10x+k ?

OpenStudy (anonymous):

you can do it but you will not get an integer

OpenStudy (anonymous):

you will get a fraction if you want a perfect square

OpenStudy (anonymous):

the way you can make sure you have a perfect square is to make sure that \[b^2-4ac=0\] so in this case you want \[(-10)^2-4\times 9\times k=0\] or \[100=36k\] \[k=\frac{100}{36}=\frac{25}{9}\]

OpenStudy (anonymous):

and then you will get \[9x^2-10x+\frac{25}{9}=\frac{1}{9}(9x-5)^2\]

OpenStudy (vanz):

Oh I see, Thank you! :)